At what temperature will the root mean square velocity of the molecule quarter that if it is at a temperature of 32 degree Celsius?

1
Expert's answer
2021-12-20T10:31:15-0500
"v_\text{rms1}=\sqrt\frac{3RT_1}{\Mu},\\\space\\
v_\text{rms2}=\frac14v_\text{rms1},\\\space\\
\sqrt\frac{3RT_2}{\Mu}=\frac14\sqrt\frac{3RT_1}{\Mu},\\\space\\
T_2=\frac{T_1}{16}=\frac{273+32}{16}=19\text{ K}."

Or -254°C.